A sample of lake water at 25C is analyzed and the following parameters are found for it: total alkalinity 6.2 10 4M phenolphthalein alkalinity 1.0 10 5 M pH 7.6 hardness 30.0 mg L1 [Mg2] 1.0 10 4M Extract all possible single ion concentrations that you can by combining one or more of these data. Also determine whether or not the water is at equilibrium with respect to the carbonate– bicarbonate system and whether or not it is saturated with calcium carbonate.
1
Expert's answer
2019-12-13T06:25:55-0500


[CO32-] = 2*10-5 M

[HCO3-] = 6.2*10-4 - 2*10-5 = 6*10-4

[H+] = 2.5*10-8 M


[Ca2+] = 1.4*10-2 M

No